OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14

OpenStudy (anonymous):

subtract 16 from both sides x+2=-2 x+2=2 solve for x

OpenStudy (anonymous):

the answered on the work don't make sense to me

OpenStudy (anonymous):

x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (anonymous):

I think it's x=0 and x=28 am I right? @KlOwNlOvE

OpenStudy (anonymous):

no x+2=2 x+2=-2 subtract 2 from both sides in the first one 2's cancel out x=0 subtract 2 by -2 both negatives so they combine x=-4

OpenStudy (anonymous):

x=-4 and x=0? @KlOwNlOvE

OpenStudy (anonymous):

yes